1+sinx /1-sinx =m^2/n^2 .then sinx is....please give me answer

Grade:10

1 Answers

manna
7 years ago
1+sinx/1-sinx+1=m^2/n^2+11+sinx /1-sinx =m^2/n^2 .then sinx is..
add1 both side
2/1-sinx=m^2/n^2+1
2/1-sinx=m^2+n^2/n^2
solve it u will get
1/2-sinx/2=n^2/m^2+n^2
sinx=1-2n^2/(m^2+n^2)